PREX/CREX Jun 4, 2019 10:00am

From PREX Wiki
Jump to: navigation, search

Back to Main Page >> PREX/CREX_Teleconferences

previous meeting << >> following meeting

Logistic information

 BlueJeans calling instructions:
 Toll-Free Number (U.S.&  Canada):  888-240-2560
 International toll number:         408-740-7256
 Bluejeans CODE:                  628 981 532
 Bluejeans link:
 Offline room:   CEBAF Center F324-325


  1. Installation progress: Kent/Ciprian Installation photo album
    • CG: installation proceeding on schedule. The vacuum problems seem to be under control.
    • CG: Collaboration buddy rule: while this is not required by the lab we would like to request everyone working after hours in the hall for the next couple of weeks make sure that they have a buddy in the hall. This person doesn't have to be working on the same thing, but both should be aware of each other and if one leaves for the day to make sure to let the other person know so that they can find an alternative buddy. People should not work by themselves in the hall after hours. KK: we should send a collaboration wide email with this. CG: will do.
  2. HRS installation update
    • DM: Left HRS leak check went really quick. Everything is back in place. AT detector will be in place and wired up in a couple of days. Main quartz to follow. Motion system is the last to check.
    • DM: the SAM work is ongoing. Reorganized the cables. We have a holder for the converters and preAmps. Two shutters installed. The software system works. We anticipate having to take the shutters out (it takes about 2h).
  3. Integrating DAQ Update
    • CC: software is in place to setup the timing for the ADC integration. This will work through GM.
    • CC: Bill Gunning put the raster back together. BM: the sync between the raster and beam is ready to go. We need to figure out if we run in beam sync or free clock (beam sync can produce density fluctuations).
  4. BPM update
    • BM: they don't need new cables run for now, we'll have to look into it further.
  5. japan progress
    • PK: several things moving forward. In particular miniruns and burpcut implementations should be ready by the start of the run. The online monitor needs some TLC. CG: will look to task Weibin with it.
  6. Injector studies:
    • CG: work ongoing and is expected to intensify in the next few days. Will be on hand this evening to help with online monitoring.
  7. Compton polarimeter
    • AZ: all is left on the photon detector is analysis of the pulser runs and check non-linearity. We have preliminary results that show below 1%. Will document the alignment and setup in log entries.
    • DG: no survey and alignment needed on the Compton. We hope to put the laser vacuum region back tomorrow and pump down Th or Fri
  8. Moller:
    • SP: Simona replaced one the PMTs for the scintillator. We will place the detector back in the box and align it.
    • SP: working on run plan. Eric is working on simulations.
  9. Beam Modulation
    • #* VO: the HW is working as expected. Working on the online plots and further analysis.
    • VO: still working on slow control processes. PK: should be ready soon to intergrate in the main CRL boot script.
  10. Pedestal studies
    • KK: 4pm will go over the schematics for the detector, BPM and BCM pedestal studies. will come up with a strategy
  11. Tune B Quad Settings
    • RyR: looked at TuneB from PREX1. Found some fluctuations and took averages to set Q2 and Q3. For Q1 we compared a couple of runs and found a scaling relation.
  12. Other:
    • SC: working through target training and will make an announcement this week
    • KK: can we get constraints on optics and main ladder? (current). Can we also get real thicknesses for these targets? SC: will work on it.


In F324

Sanghwa P, KK, Ciprian G, Tao Y, Marrisa P, Catherine ?, Yufan Chen, Cameron C, Paul K, Michael , Robert R, Devika P,


Adam Z, Chandan G, Dave G, Devi A, Eric K, Juliette M, Quinn Campagna, Ryan R, Silviu D, Siyu J, Richard H, Amali P, Bob M, Alexandre C, Dustin M, Victoria O,

Cannot make it

Kent P, Caryn P,